OpenStudy (anonymous):

Solve 3x^2 = -12x - 15

OpenStudy (anonymous):

@esamalaa

OpenStudy (triciaal):

rearrange to set =0 then factor

OpenStudy (anonymous):

I don't get it.

OpenStudy (triciaal):

what do you have to do to this equation to let the value on the right = 0

OpenStudy (anonymous):

@esamalaa

OpenStudy (anonymous):

subtract 3x^2?

OpenStudy (triciaal):

what you do to one side you also do to the other to keep the balance

OpenStudy (anonymous):

im confused..

OpenStudy (triciaal):

Solve 3x^2 = -12x - 15

OpenStudy (anonymous):

-2(p or m)1i

OpenStudy (triciaal):

so if you subtract from the left you have to subtract from the right as well

OpenStudy (anonymous):

x=−2+i,−2−i

OpenStudy (triciaal):

3x^2 + 12 x + 15 = 0 we can divide everything by 3 3(x^2 + 4x +5)=0